# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manusript

### v4

@@ I Introduction
A possible solution to the above limitations, which has proven advantageous in various ML fields, is the utilization of unlabeled data [11, 12]. By training models on a large corpus of unlabeled data via self-supervised learning (SSL), the feature extractor is able to learn general representations and avoid overfitting on its training distributions. Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on either electrocardiography, electromyography or electroencephalography signals (see Section II-B). Therefore, in this work we turn our attention to the Phonocardiogram domain and propose training robust classifiers for PCG signal classification via contrastive self-supervised learning, leveraging both labeled and unlabeled datasets. Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13]. The design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework. To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations () and models () applied on PCG signals and report their effect on training and generalization. Specifically, we attempt to provide insight into the following research questions:
Does the application of contrastive SSL prove a valid training method for extracting robust and meaningful representations in the PCG classification domain?
@@
To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets; (b) examine the effectiveness on a global “normal vs. abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable); (c) evaluate all trained classifiers on OOD data that are left out during training; and finally, (d) calculate the effect size of each augmentation. To the best of our knowledge, this is the first paper to research the effect of each applied augmentation on learning robust OOD PCG representations. To facilitate further research in the field and provide a comprehensive baseline, we have open-sourced our codebase111Code available at: https://github.com/aristotelisballas/listen2yourheart.
@@ II-B Self-Supervised Learning in 1D biosignal classification
Self-supervised representation learning (SSL) [28] was initially proposed for mitigating the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks [29] to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding [30], the goal of Self-Supervised Contrastive Learning [31, 32] is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions [33]. The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which capable of leveraging small amounts of data while yielding adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [34]. The authors of [35] propose augmenting multi-channel EEG signals by recombining several channels into additional ones which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying predefined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of [36] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in [37] where contrastive predictive coding is implemented on similar SSL-learned features to predict sleep stages and detect pathologies in multi-channel EEG signals. Finally, [38] proposed BENDR, which combines attention mechanisms [39] and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well [40], most of which focus on either emotion recognition or pathology detection. For example, in [41] the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the-art results. In another paper studying emotion recognition from ECG signals, [42] introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44] or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, [46] propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models [47]. For example, [48] introduces NeuroPose, a system for 3D hand pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
@@ II-C Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [50]. To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data-generating processes. Here, we reference prior works which fall under the above research fields.
@@
Finally, Domain Generalization (DG) algorithms aim to tackle model generalization at its core, as the main goal is to provide models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of [62] propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, [63] and [64] propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of [4] introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from [65], for 12-lead ECG and 64-channel EEG classification.
@@ III Materials and Methods
In this work, we introduce an evaluation framework for assessing the combination of various transformations, as well as the robustness of a network trained via contrastive SSL, against a fully supervised baseline model. This allows us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain, while also highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
@@ III-A Datasets & Downstream tasks
The Ephnogram [67] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals making it appropriate for developing noise-resilient methods. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled PCG signals for the SSL pretraining of our models. All recordings were sampled at 8 kHz.
@@
The Pascal [69] Heart Sound Classification Challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole222Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced as publicly available challenges for PCG classification. Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart sound recordings as a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 Challenge [15] was designed for the development of algorithms that can detect heart murmur in PCG signals. Building on the 2016 challenge, the 2022 edition contains a diverse dataset of heart sound recordings collected from multiple auscultation locations. The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals recorded from one of the four heart valves or another prominent auscultation location. Each recording is classified into one of three categories: present (heart murmur detected), absent (no heart murmur detected), or unknown (e.g., due to noise or corrupted audio), rendering this a multi-class classification task. All recordings were sampled at a rate of 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process to [27] for converting all signals into a common format. As a first step and to avoid transient noise artifacts, we discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart sound recording333In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage we assign pseudolabels to each window, as described in Section III-B. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can also be evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section III-D.
@@ III-B Self-Supervised Contrastive Learning for PCG Representation Learning
For the contrastive prediction task, we employ the NT-Xent loss function introduced in [30] and implemented in the SimCLR framework [70]. Specifically, given a set which includes an augmented positive pair and , a model trained via contrastive SSL aims to identify in for a given . The contrastive loss for a positive pair is formulated as follows:
where is the initial batch size of the randomly sampled signals, is an indicator function and a temperature parameter. The function calculates the cosine similarity between the encoded vectors and is defined as:
The final loss is computed across all positive pairs in a mini-batch.
@@ III-C Augmentations for PCG Representation Learning
Arguably, the most integral part of the proposed framework is the family of transformation functions , which is applied on each audio window, as the type and intensity of each function can drastically affect the learning of the encoder network. In order to conduct a rigorous evaluation of all augmentations included in the proposed contrastive SSL framework and also report the functions which should most likely be applied for robust PCG representation learning, we select to implement functions that either simulate factors that can be found in real world scenarios. For example, the first thing that comes to mind is the volume of noise polluting recordings during PCG collection. Additionally, the difference in sensor conductance can lead to different scales of the recorded signals. To this end, it would make sense to at least either add artificial noise to randomly scale each audio window during the pretraining task. Following the above rationale, in our study we select to implement and evaluate the effectiveness of a total of 6 different augmentations, each of which is described below:
@@
Reverse: The signal sequence is reversed.
@@ III-D Evaluation Protocol
In the first step of the framework, we create a pool of 5 sec unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and Fpcgdb. After the preprocessing step has been completed and all signals are homogenized in the common format described in Section III-A, they are randomly shuffled and split into training and validation splits (80% and 20% respectively). In the second step, the split signals are cut into batches. Each sample in the batch is then duplicated and augmented by applying either one or a combination of the functions described in Section III-C. The augmented samples are then passed through a Neural Network encoder, which is trained via the contrastive pretraining task (Step 3). After the SSL backbone model has been trained, its weights are frozen.

### v5

# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manuscript

@@ I Introduction
A possible solution to the above limitations, which has proven advantageous in various ML fields, is the utilization of unlabeled data [11, 12]. By training models on a large corpus of unlabeled data via self-supervised learning (SSL), the feature extractor is able to learn general representations and avoid overfitting on its training distributions. Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on electrocardiography, electromyography or electroencephalography signals (see Section II-B). Therefore, in this work we turn our attention to the Phonocardiogram domain and propose training robust classifiers for PCG signal classification via contrastive self-supervised learning, leveraging both labeled and unlabeled datasets. Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13], the design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework. To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations (~) and models (~) applied on PCG signals and report their effect on training and generalization. Specifically, we attempt to provide insight into the following research questions:
Does the application of contrastive SSL prove to be a valid training method for extracting robust and meaningful representations in the PCG classification domain?
@@
To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets, (b) examine the effectiveness on a global “normal vs. abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable), (c) evaluate all trained classifiers on OOD data that are left out during training, and finally, (d) calculate the effect size of each augmentation. To the best of our knowledge, this is the first paper to research the effect of each applied augmentation on learning robust OOD PCG representations. To facilitate further research in the field and provide a comprehensive baseline, we have open-sourced our codebase111Code available at: https://github.com/aristotelisballas/listen2yourheart.
@@ II-B Self-Supervised Learning in 1D biosignal classification
Self-supervised representation learning (SSL) [28] was initially proposed for mitigating the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks [29] to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding [30], the goal of Self-Supervised Contrastive Learning [31, 32] is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions [33]. The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there is a need to provide systems which are capable of leveraging small amounts of data while yielding adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG), or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [34]. The authors of [35] propose augmenting multi-channel EEG signals by recombining several channels into additional ones which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying predefined transformations and employ contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of [36] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in [37] where contrastive predictive coding is implemented on similar SSL-learned features to predict sleep stages and detect pathologies in multi-channel EEG signals. Finally, [38] proposed BENDR, which combines attention mechanisms [39] and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well [40], most of which focus on either emotion recognition or pathology detection. For example, in [41] the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the-art results. Another paper [42] studying emotion recognition from ECG signals introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44], or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, [46] propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field that can greatly benefit from the application of robust prediction models [47]. For example, [48] introduces NeuroPose, a system for 3D hand-pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a another user. Similarly, the authors of [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
@@ II-C Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [50]. To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-IID distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data-generating processes. Here, we reference prior works which fall under the above research fields.
@@
Finally, Domain Generalization (DG) algorithms aim to tackle model generalization at its core, as the main goal is to provide models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of [62] propose an ensemble classifier fusion method that yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, [63] and [64] propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of [4] introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from [65], for 12-lead ECG and 64-channel EEG classification.
@@ III Materials and Methods
In this work, we introduce an evaluation framework for assessing the combination of various transformations, as well as the robustness of a network trained via contrastive SSL, against a fully supervised baseline model. This enables us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain as well as highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
@@ III-A Datasets & Downstream tasks
The Ephnogram [67] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30 minutes long while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor-quality sample signals making it appropriate for developing noise-resilient methods. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled PCG signals for the SSL pretraining of our models. All recordings were sampled at 8 kHz.
@@
The Pascal [69] Heart Sound Classification Challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenge contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole222Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced as publicly available challenges for PCG classification. Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart-sound recordings as a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 Challenge [15] was designed for the development of algorithms that can detect heart murmur in PCG signals. Building on the 2016 challenge, the 2022 edition contains a diverse dataset of heart sound recordings collected from multiple auscultation locations. The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals, recorded from one of the four heart valves or another prominent auscultation location. Each recording is classified into one of three categories: present (heart murmur detected), absent (no heart murmur detected), or unknown (e.g., due to noise or corrupted audio), rendering this a multi-class classification task. All recordings were sampled at a rate of 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process to [27] for converting all signals into a common format. As a first step and to avoid transient noise artifacts, we discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5-second, overlapping windows with 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart-sound recording333In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage we assign pseudolabels to each window, as described in Section III-B. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can also be evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section III-D.
@@ III-B Self-Supervised Contrastive Learning for PCG Representation Learning
For the contrastive prediction task, we employ the NT-Xent loss function introduced in [30] and implemented in the SimCLR framework [70]. Specifically, given a set which includes an augmented sample pair and , a model trained via contrastive SSL aims to identify in for a given . The contrastive loss applied on the representations of any pair is formulated as follows:
where and are the embedded representations of and respectively, is the number of the initially randomly sampled signals (before the augmentations, i.e., ), is an indicator function and a temperature parameter. The function calculates the cosine similarity between the encoded vectors and is defined as:
The final loss is computed across all pairs in a mini-batch.
@@ III-C Augmentations for PCG Representation Learning
Arguably, the most integral part of the proposed framework is the family of transformation functions , which is applied on each audio window, as the type and intensity of each function can drastically affect the learning of the encoder network. In order to conduct a rigorous evaluation of all augmentations included in the proposed contrastive SSL framework and also report the functions which should most likely be applied for robust PCG representation learning, we select to implement functions that either simulate factors that can be found in real world scenarios, such as the level (volume) of noise. Additionally, the difference in sensor conductance can lead to different scales of the recorded signals. To this end, it would make sense to at least either add artificial noise to randomly scale each audio window during the pretraining task. Following the above rationale, in our study we select to implement and evaluate the effectiveness of a total of 6 different augmentations, each of which is described below:
@@
Reverse: The signal sequence is reversed, i.e., where is the length of signal .
@@ III-D Evaluation Protocol
In the first step of the framework, we create a pool of 5-second, unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and FPCGDB. After the preprocessing step has been completed and all signals are homogenized in the common format described in Section III-A, they are randomly shuffled and split into training and validation splits (80% and 20% respectively). In the second step, the split signals are cut into batches. Each sample in the batch is then duplicated and augmented by applying either one or a combination of the functions described in Section III-C. The augmented samples are then passed through a Neural Network encoder, which is trained via the contrastive pretraining task (Step 3). After the SSL backbone model has been trained, its weights are frozen.